# Accretion flow around Kerr metric in the infra-red limit of asymptotically safe gravity

### v1

@@ IV Quantum-Correction Effects on QPOs: Power Spectral Density Analysis
Fig.6 presents the PSD analysis computed from the temporal oscillations of the mass accretion rate after the system has reached a steady-state, for the Schwarzschild spacetime and the quantum corrected IR1 and IR2 models. The PSD at $r=2.3M$ (left panel) and at $r=6.11M$ (right panel), corresponding to the region near the horizon and the vicinity of the ISCO, respectively. We found similar characteristic frequencies are observed at both radii therby the detected QPOs correspond to global oscillation modes of the accretion flow rather than arising from local turbulence or numerical artifacts. In other words, the origin of the QPOs in these systems is physical and can be attributed to the excitation of advective-acoustic modes driven by the shock cone formed around the black hole.
@@
A physically motivated explanation for this behaviour is the strong frame-dragging induced by rapid rotation, which reorganises the inner accretion flow and shifts the most efficient modulation region outward. Near the black-hole horizon, the mass flux through a small extraction sphere becomes comparatively steady, or dominated by incoherent high-shear variability, leading to suppressed PSD amplitudes. While, near the ISCO the shock cone breathing and post-shock recirculation remain dynamically active, periodically loading and unloading the accretion flow and producing stronger quasi-periodic modulation.
@@
V Identification of Observable QPO Frequencies from Numerical PSDs
In this section, we investigate the physical validity of the QPO frequencies obtained numerically from the PSD analysis and discuss their consistency with astrophysical observations. Throughout this work, in order to express the numerically calculated frequencies in the SI unit system, we adopt a fiducial stellar-mass black hole with mass $M=10\,M_{\odot}$. Consequently, the QPO frequencies identified in the PSD analyses are directly comparable to those observed in Galactic black hole X-ray binaries (XRBs) Remillard and McClintock (2006b). On the other hand, these frequencies can be straightforwardly rescaled, allowing predictions and comparisons to be made for systems hosting massive and supermassive black holes.
As discussed in detail in Section IV, for both non-rotating and rotating black holes in the infra-red limit of asymptotically safe gravity, the QPO frequencies exhibit systematic variations and the emergence of overtones depending on the black hole spin parameter and the quantum correction term. Here, we explicitly demonstrate how these numerical results are consistent with observations. Since the frequencies identified in Section IV are obtained by adopting $M=10\,M_{\odot}$, the resulting QPO frequencies naturally fall within the range of low- and moderately high-frequency QPOs observed in Galactic XRBs Donmez (2025).
@@ VI Conclusion
Global oscillation modes are trapped and sustained by the shock cone produced, as evidenced by the time-dependent behaviour of the mass accretion rate. The ensuing QPOs can be identified through further PSD analysis. The tests demonstrate that the QPO frequencies are identical regardless of the extraction radius when PSDs are computed at two distinct radial locations. This indicates that the observed QPOs are real physical signals originating from global oscillation modes of the accretion flow rather than numerical errors. The simulations show the existence of basic modes, harmonic overtones, and near-commensurate frequency ratios like 2:1 and 3:2, depending on the black hole spin and the quantum correction parameters. It is found that totally nonlinear advective-acoustic cycles functioning inside the post-shock cavity are the source of these ratios. The numerical results further the creation of stronger and more distinct oscillation modes is facilitated by modest quantum corrections and moderate black hole rotation. Large quantum correction values, on the other hand, diminish the efficiency with which the post-shock cavity sustains prolonged oscillations by weakening shock compression and suppressing oscillation amplitudes.
@@
Using a stellar-mass black hole $M=10\,M_{\odot}$, the numerically calculated QPO frequencies naturally fall within the range of low- and moderately high-frequency QPOs observed in XRBs. The numerical frequencies are consistent with those observed in Type-C QPO systems. The harmonic structures and near 3:23\!:\!2 frequency ratios in the simulations are in close agreement with the features reported in observational timing studies. This demonstrates that shock-driven global oscillation modes alone can reproduce many of the fundamental QPO characteristics without invoking epicyclic resonance mechanisms associated with thin accretion disks. On the other hand, the inverse mass-scaling relation shows that the frequencies and physical mechanisms identified for XRBs are also applicable to massive black holes. Using this scaling, QPO frequencies in the milli-Hertz range are predicted for massive black holes, in agreement with observational results reported in the literature Gierlinski et al. (2008); Ren et al. (2023); Reis et al. (2012). These findings thus demonstrate QPOs as a physically reliable and mass-scalable probe of strong-field gravity, offering an additional observational channel to test infrared-modified spacetime geometries over the whole black hole mass spectrum.

### v2

@@ 4 Quantum-Correction Effects on QPOs: Power Spectral Density Analysis
Fig.6 presents the PSD analysis computed from the temporal oscillations of the mass accretion rate after the system has reached a steady-state, for the Schwarzschild spacetime and the quantum corrected IR1 and IR2 models. The PSD in $r=2.3M$ (left panel) and in $r=6.11M$ (right panel), corresponding to the region near the horizon and a larger-radius extraction region, respectively. We found similar characteristic frequencies are observed at both radii thereby the detected QPOs correspond to global oscillation modes of the accretion flow rather than arising from local turbulence or numerical artifacts. In other words, the origin of the QPOs in these systems is physical and can be attributed to the excitation of advective-acoustic modes driven by the shock cone formed around the black hole.
@@
A physically motivated explanation for this behaviour is the strong frame-dragging induced by rapid rotation, which reorganises the inner accretion flow and shifts the most efficient modulation region outward. Near the black-hole horizon, the mass flux through a small extraction sphere becomes comparatively steady, or dominated by incoherent high-shear variability, leading to suppressed PSD amplitudes. While, at larger radii the shock cone breathing and post-shock recirculation remain dynamically active, periodically loading and unloading the accretion flow and producing stronger quasi-periodic modulation.
@@
5 Identification of Observable QPO Frequencies from Numerical PSDs
In this section, we investigate the physical validity of the QPO frequencies obtained numerically from the PSD analysis and discuss their consistency with astrophysical observations. Throughout this work, in order to express the numerically calculated frequencies in the SI unit system, we adopt a fiducial stellar-mass black hole with mass $M=10\,M_{\odot}$. Consequently, the QPO frequencies identified in the PSD analyses are directly comparable to those observed in Galactic black hole X-ray binaries (XRBs) Remillard and McClintock (2006b). However, these frequencies can be straightforwardly rescaled, allowing predictions and comparisons to be made for systems hosting massive and supermassive black holes.
As discussed in detail in Section 4, for both non-rotating and rotating black holes in the infra-red limit of asymptotically safe gravity, the QPO frequencies exhibit systematic variations and the emergence of overtones depending on the black hole spin parameter and the quantum correction term. Here, we explicitly demonstrate how these numerical results are consistent with observations. Since the frequencies identified in Section 4 are obtained by adopting $M=10\,M_{\odot}$, the resulting QPO frequencies naturally fall within the range of low- and moderately high-frequency QPOs observed in Galactic XRBs Donmez (2025).
@@ 5 Identification of Observable QPO Frequencies from Numerical PSDs
From an observational perspective, our numerical results indicate that variations in the black hole spin, accretion state, and the initial flow parameters of the BHL mechanism may introduce partial degeneracies with the effects induced by the quantum-correction parameter $\xi$. Nevertheless, our findings show that increasing $\xi$ leads to a systematic suppression of QPO peaks in the PSD analysis, a reduction in harmonic richness, and a damping of the coherence of oscillation modes. In contrast, variations in the black hole spin primarily modify the azimuthal dynamics of the accreting matter and the overall efficiency of flow modulation. These differences are reflected in distinct trends of the fractional RMS variability shown in the bottom-right panel of Fig.5. At the same time, the presence and relative strength of harmonic peaks, as well as the coherence properties of the QPOs, further highlight these differences. Such signatures correspond to observable physical variations, indicating that shock-driven timing diagnostics provide an independent channel for probing infrared-modified gravity effects.
@@ 6 Conclusion
Global oscillation modes are trapped and sustained within the post-shock region, as evidenced by the time-dependent behaviour of the mass accretion rate. The associated QPOs are identified through power spectral density (PSD) analyses performed after the system has reached a quasi-steady state. Importantly, identical characteristic frequencies are obtained when the PSDs are computed at two physically distinct extraction radii, namely near the compact object and at a larger extraction radius. This radius-independent behaviour demonstrates that the detected QPOs correspond to coherent global oscillation modes of the accretion flow and are not artifacts of a particular sampling location or transient numerical effects. We further note that the global nature of the modes produced by the same numerical framework was previously established through a dedicated time-window analysis (see Fig. 2 of Ref. Donmez and Dogan (2024)), where the persistence of the characteristic frequencies over different time intervals was demonstrated. The simulations show the existence of basic modes, harmonic overtones, and near-commensurate frequency ratios like 2:1 and 3:2, depending on the black hole spin and the quantum correction parameters. It is found that totally nonlinear advective-acoustic cycles functioning inside the post-shock cavity are the source of these ratios. The numerical results further the creation of stronger and more distinct oscillation modes is facilitated by modest quantum corrections and moderate black hole rotation. Large quantum correction values, on the other hand, diminish the efficiency with which the post-shock cavity sustains prolonged oscillations by weakening shock compression and suppressing oscillation amplitudes.
@@
Using a stellar-mass black hole $M=10\,M_{\odot}$, the numerically calculated QPO frequencies naturally fall within the range of low- and moderately high-frequency QPOs observed in XRBs. The numerical frequencies fall within the characteristic frequency range of Type-C QPOs observed in Galactic X-ray binaries and reproduce several of their key phenomenological features, such as strong coherence and harmonic structure, when scaled to a stellar-mass black hole. The harmonic structures and near 3:23\!:\!2 frequency ratios in the simulations are in close agreement with the features reported in observational timing studies. This demonstrates that Shock-driven global oscillation modes can reproduce many of the fundamental QPO characteristics without explicitly invoking epicyclic resonance mechanisms, providing a complementary physical explanation to disk-based models. On the other hand, the inverse mass-scaling relation shows that the frequencies and physical mechanisms identified for XRBs are also applicable to massive black holes. Using this scaling, QPO frequencies in the milli-Hertz range are predicted for massive black holes, in agreement with observational results reported in the literature Gierlinski et al. (2008); Ren et al. (2023); Reis et al. (2012). These findings thus demonstrate QPOs as a physically reliable and mass-scalable probe of strong-field gravity, offering an additional observational channel to test infrared-modified spacetime geometries over the whole black hole mass spectrum.
A full observational disentanglement of black hole spin, accretion conditions, and infrared quantum-gravity effects will ultimately require a multi-observable approach that combines timing diagnostics with spectral, imaging, and independent spin measurements. In contrast, in the present study we focus on shock-driven variability and QPO properties obtained from relativistic hydrodynamic simulations, while recognizing that isolating the impact of the quantum-correction parameter will require breaking parameter degeneracies through complementary observational techniques. In this context, the time-dependent signatures identified in this work—such as variations in QPO peak amplitudes, harmonic structures, coherence properties, and fractional RMS variability that represent independent and physically motivated numerical results that probe the strong-gravity regime. In future studies, we aim to construct a systematic framework by incorporating these observables and exploring a broader parameter space in order to assess their combined diagnostic power. Such an approach will enable a more detailed understanding of the physical consequences of infrared-modified spacetime properties around black holes and provide clearer guidance on how these effects may be detected observationally.